# Topological bounds on the dynamical growth rate of chemical reaction networks

### 03 Mar 2026 v1

@@ B. Maximum Amplification Factor (MAF) results
For solver implementations of the above, see [6, 1].
@@
Von Neumann in [5] and Theorem 9.8 in [4] rigorously prove that $0<\alpha<\infty$ for semi-autonomous networks. Since autonomous networks are also semi-autonomous, the bounds trivially extend to them as well. While detailed proofs of the statement can be found there, we sketch a heuristic derivation of the same.
@@
Von Neumann in [5] and Theorem 9.10 in [4] prove that there exists a unique MAF for irreducible semi-autonomous networks. Since autonomous networks are also semi-autonomous, irreducible autonomous networks have a unique MAF.
@@
with the constraint $\underline{\mathbf{x}}>\mathbf{0}$. The relation between solutions of Eq. (S7) and the generalized eigenvalue problem is studied in Ref. [7].
An autonomous network is minimal if there is no subnetwork that is also autonomous. Since there is no species that is never consumed, such a network must also be irreducible. It has been shown that minimal autonomous networks have an equal number of species and reactions [3, 2, 8], and that their input matrix can be rearranged into a diagonal form. Clearly, then, the network is irreducible and $\mathbb{S}^{-}$ is non-singular. Using the results in [7], the problem of finding MAF reduces to the generalized eigenvalue problem of Eq. (S10), which further reduces to the standard eigenvalue problem

### 30 Mar 2026 v2

@@ B. Maximum Amplification Factor (MAF) results
For solver implementations of the above, see [sargent2019neumann, blanco2025identifyingselfamplifyinghypergraphstructures].
@@
Von Neumann in [neumann1945model] and Theorem 9.8 in [gale1989theory] rigorously prove that $0<\alpha<\infty$ for semi-autonomous networks. Since autonomous networks are also semi-autonomous, the bounds trivially extend to them as well. While detailed proofs of the statement can be found there, we sketch a heuristic derivation of the same.
@@
Von Neumann in [neumann1945model] and Theorem 9.10 in [gale1989theory] prove that there exists a unique MAF for irreducible semi-autonomous networks. Since autonomous networks are also semi-autonomous, irreducible autonomous networks have a unique MAF.
@@
with the constraint $\underline{\mathbf{x}}>\mathbf{0}$. The relation between solutions of Eq. (S7) and the generalized eigenvalue problem is studied in Ref. [thompson1971neumann].
An autonomous network is minimal if there is no subnetwork that is also autonomous. Since there is no species that is never consumed, such a network must also be irreducible. It has been shown that minimal autonomous networks have an equal number of species and reactions [gagrani2024polyhedral, blokhuis2020universal, vassena2024unstable], and that their input matrix can be rearranged into a diagonal form. Clearly, then, the network is irreducible and $\mathbb{S}^{-}$ is non-singular. Using the results in [thompson1971neumann], the problem of finding MAF reduces to the generalized eigenvalue problem of Eq. (S10), which further reduces to the standard eigenvalue problem